OpenStudy (nerdgamer_):

Which mixed number falls between four and three-eighths and five and five-twelfths? (A) four and one-fourth (B) five and two-thirds (C) four and one-half (D) five and five-sixths

OpenStudy (nerdgamer_):

Medal and A fan to whoever help me!!! :)

OpenStudy (anonymous):

All of the fractions can be written with a denominator of 24

OpenStudy (nerdgamer_):

ok :3

OpenStudy (nerdgamer_):

keep going ........

OpenStudy (nerdgamer_):

Im in a hurry make it fast please:3

OpenStudy (anonymous):

Or you could put them as a decimal and then put them in order like that.

OpenStudy (nerdgamer_):

ok How do I get the answer?

OpenStudy (anonymous):

4 3/8 = 4.375 and 5 5/12= 5.416 A. 4 1/4 = 4.25 B. 5 2/3 = 5.6 C. 4 1/2 = 4.5 D. 5 5/6 = 5.8

OpenStudy (anonymous):

On a number line it would look like this 4.25, 4.37, 4.5, 5.4, 5.6

OpenStudy (nerdgamer_):

okayyyy... and

OpenStudy (anonymous):

So what is the answer?

OpenStudy (nerdgamer_):

is it ummm... (C) ???

OpenStudy (anonymous):

Yes! 4.5 would be the correct answer :)

OpenStudy (nerdgamer_):

Yahhhyyy thxs!! soo much bye

OpenStudy (anonymous):

Yw